State the hypotheses and identify the claim, find the critical value(s), compute the test value, make the decision, summarize the results (make the appropriate statement of the results of the claim included rejection or non-rejection of the null hypothesis). Please show all of your work.

A researcher wishes to see if students show a time preference for anatomy labs. A sample of four lab classes shows enrollment.

At = 0.10, do the students show a time preference for the labs?

8AM 12PM 2PM 4PM

Students 27 33 38 24
Answer #1

Calculate the expected value. 27+33+38+24 E= 4 122 4 = 30.5 Null hypothesis, H: the students do not shows time preferences fo

Calculate the chi-square test statistic value. (0-E) χ = Σ- E = 0.402+0.205 +1.844+1.385 = 3.836 Calculate the degrees of fre